Graduate high school students who passed, 125. Undergraduate students 280. Total pass,405.
Graduates who failed 55, Undergraduates who failed were 50. Total failed together, both graduates and undergraduates were 105, and total passed together were 405. How would I solve if given this problem?
a. P(Pass)
b. P(Pass and Graduate)
c.P(Fail/undergraduate /Fail)
D.P(Graduate/Fail)
and
e. P(Pass or Undergraduate)

Answer by stanbon(75887)

Graduate high school students who passed, 125.
Undergraduate students who passed, 280
Total pass,405.
------------------------------------
Graduates who failed 55,
Undergraduates who failed were 50.
Total failed together, both graduates and undergraduates were 105
----------------------
Grand total:: 405+105 = 510
-----------------------------------
Keep in mind::
P(a AND b) = P(a)+P(b)-P(a OR b)
P(a | b) = P(a AND b)/P(b)
---------------------------
How would I solve if given this problem?
a. P(Pass) = 405/510
---------------------------------
b. P(Pass AND Graduate) = 125/510
-------------------------------------------------
c.P(Fail/undergraduate) = 50/510
-------------------------------------
D.P(Graduate/Fail) = 55/105
-------------------------------------
e. P(Pass or Undergraduate) = (405+(280+50) - 280)/510 = 455/510
---------------------